# Yong Ji1, Junye Li1 and Rui Yang∗2,3,4

### v2

@@
Abstract. Saturated set and its reduced case, the set of generic points, constitute two significant types of fractal-like sets in multifractal analysis of dynamical systems. In the context of infinite entropy systems, this paper aims to give some qualitative aspects of saturated sets and the set of generic points in both topological and measure-theoretic perspectives. For systems with specification property, we establish the certain variational principles for saturated sets in terms of Bowen and packing metric mean dimensions, and show the upper capacity metric mean dimension of saturated sets have full metric mean dimension. All results are useful for understanding the topological structures of dynamical systems with infinite topological entropy.
As applications, we further exhibit some qualitative aspects of metric mean dimensions of level sets and the set of mean Li-Yorke pairs in infinite entropy systems.
@@ 1 Introduction
Let $(X,d,f)$ be a topological dynamical system (t.d.s. for short), where $(X,d)$ is a compact metric space with a metric $d$, and $f:X\to X$ is a homeomorphism. Let $M(X)$, $M_{f}(X)$, $M_{f}^{e}(x)$ denote the sets of Borel probability measures endowed with weak-topology, $f$-invariant Borel probability measures, and ergodic $f$-invariant Borel probability measures on $X$, respectively. Measure-theoretic entropy and topological entropy are two vital invariants for us to understand the topological complexity of dynamical systems, which are related by the classical variational principle:
@@
where $h_{top}(f)$ is the topological entropy of $(X,f)$, $h_{\mu}(f)$ is the measure-theoretic entropy of $\mu$, and the supremum ranges over all $f$-invariant measures of $X$. In 1973, analogous to the definition of Hausdorff dimension Bowen [Bow73] introduced Bowen topological entropy for any subset of the phase space. The pioneering work directly contributes to the development of dimension theory of dynamical systems, especially as a powerful tool to study the “size” of fractal-like sets in multifractal analysis.
The multifractal analysis, which can date back to Besicovitch, takes into account the decomposition of the whole phase space into smaller subsets consisting of the points with the similar dynamical behaviors(or known as multifractal spectrum), and then from the geometrical or topological viewpoint describes the “size” of those subsets by using fractal dimensions and dimensional entropies. Such a procedure allows us to get more information about the dynamics and even recovers the partial dynamical behaviors of the system, which is called phenomenon multifractal rigidity. We aim to study the multifractal analysis of saturated sets in infinite entropy systems.
Let us mention some relevant results about the dimensional entropies of saturated sets. Given $x\in X$, by $\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}\in M(X)$, where $\delta_{x}$ is the Dirac mass at $x$, we denote the $n$-th empirical measure of the Birkhoff average w.r.t. $x$. Then the limit-point set $V_{f}(x)$ of $\{\mathcal{E}_{n}(x)\}_{n\geq 1}$ is a nonempty compact connected subset of $M_{f}(X)$ in the sense of weak-topology (cf. [DGS76, Proposition 3.8]). Specially, when $V_{f}(x)=\{\mu\}$ for some measure $\mu$, one says that $x$ is a generic point of $\mu$. The set of all generic points of $\mu$ is denoted by $G_{\mu}$. Then the Birkhoff’s ergodic theorem implies that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic. Bowen [Bow73] proved that for any ergodic measure $\mu$, the Bowen topological entropy of $G_{\mu}$ is equal to $h_{\mu}(f)$, that is,
@@
For non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$ and even easily constructs the example showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$ such that Bowen’s result is not valid. It is natural to expect whether there are some proper conditions such that the equality still holds for non-ergodic measures. A subset $D$ of $X$ is called a saturated set if $x\in D$ and $V_{f}(x)=V_{f}(y)$, then $y\in D$. More generally, given a nonempty closed connected subset $K\subset M_{f}(X)$, we can obtain a saturated set $G_{K}=\{x\in X:V_{f}(x)=K\}$ associated with $K$. For systems with specification property, including hyperbolic systems, Sigmund [Sig72, Sig74] firstly proved the existence of the saturated sets, which was later extended to non-uniformly hyperbolic and non-uniformly expanding systems by several authors [LST13, TV17]. As for the estimation of Bowen topological entropy of saturated sets, for systems with $g$-almost product property that is weaker than specification property, which solely requires a point partially $\epsilon$-shadowing the given finite pieces of orbit-segments, and with uniform separation property (e.g. expansive and more generally asymptotically $h$-expansive systems), Pfister and Sullivan [PS07] established the certain variational principle for any nonempty compact connected set $K\subset M_{f}(X)$,
@@
Specially, only the assumption of $g$-almost product property they [PS07] showed
$h_{top}^{B}(f,G_{\mu})=h_{\mu}(f)$
holds for all $\mu\in M_{f}(X)$, which extends Bowen’s remarkable result to non-ergodic measures. Consequently, the work of [Bow73, PS07] leads to an intensive study toward the estimation of different types of dimensional entropies of saturated sets. Packing entropy of non-compact subsets was introduced by Feng and Huang [FH12] in a way resembling the definition of packing dimension. Additionally, under the (stronger) conditions of specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] showed the following variational principle for any nonempty compact connected set $K\subset M_{f}(X)$,
$h_{top}^{P}(f,G_{K})=\sup\{h_{\mu}(f):\mu\in K\},$
where $h_{top}^{P}(f,K)$ is the packing entropy of $K$. Afterwards, weakening the specification property, for systems with $g$-almost product property, Hou, Tian and Zhang [HTZ23] not only proved that the variational principle $h_{top}^{P}(f,G_{K}^{C})=\sup\{h_{\mu}(f):\mu\in K\}$ holds for (mixed) saturated sets with more complicated dynamics, but also showed the upper capacity topological entropy of the saturated sets have full topological entropy, i.e.
@@
where $G_{K}^{C}:=G_{K}\cap\{x\in X:C_{f}(X)\subset\omega_{f}(x)\},$ $\omega_{f}(x):=\bigcap_{n=0}^{\infty}\overline{\bigcup_{k=n}^{\infty}\{f^{k}x\}}$ is the $\omega$-limit set of $x$, and $C_{f}(X)$ is the measure center given by $C_{f}(X):=\overline{\cup_{\mu\in M_{f}(X)}{\rm supp}{(\mu)}}$, ${\rm supp}(\mu)$ is the supported set of $\mu$. Furthermore, for systems satisfying $g$-almost product property and uniform separation property, Huang, Tian and Wang [HTW19] showed Pfister and Sullivan’s variational principle is valid for transitive-saturated set $G_{K}^{T}:=G_{K}\cap{\rm Tran}$, where ${\rm Tran}$ denotes the set of transitive points. Readers can turn to [JCZ22, LLS17] for more results involving the dimensional entropies of saturated sets in the setting of nonuniformly hyperbolic systems, and [ZC18, DZZ23, Z18, RTZ23, ZRZ23] for some results concerning the amenable Bowen and packing topological entropies of saturated sets in the context of the actions of amenable groups.
According to the possible values of topological entropy, dynamical systems can be divided into three types: zero entropy systems, finite positive entropy systems and infinite entropy systems. Infinite entropy system always means the system may possess quiet complicated topological complexity, so the notion of topological entropy no longer provides more information about the dynamics of the system. In 2000, Lindenstrauss and Weiss [LW00] introduced a refined notion for topological entropy, called metric mean dimension, which is a dynamical analogue of box dimension to measure how fast the $\epsilon$-topological entropy diverges to $\infty$ as $\epsilon$ goes to 00. It turns out that any finite entropy system has zero metric mean dimension. See some applications of metric mean dimension in the analog compression of information theory [GS20], and in estimating the upper bound of mean dimension of some complex systems [Tsu18a, Tsu18b]. The notion of measure-theoretic counterpart of metric mean dimension, called upper and lower mean Rényi information dimensions, is due to Gutman and Śpiewak[GS21]. More precisely, the lower and upper mean Rényi information dimensions of $\mu\in M_{f}(X)$ are respectively defined by
@@
which can be explained how fast the Kolmogorov-Sinai $\epsilon$-entropy $\inf_{\mathrm{diam}\xi<\epsilon}\limits h_{\mu}(f,\xi)$ diverges to $\infty$ as $\epsilon\to 0$. Some variational principles of metric mean dimension of $\mathbb{Z}$-actions can be found in [LW00, GS21, S22, YCZ23a]. Then the aforementioned work suggests the following questions for saturated sets in infinite entropy systems:
@@
    For ergodic measures, do the metric mean dimensions of the set of generic points equal to the mean Rényi information dimensions?
@@
    For non-ergodic measures, under the assumption of specification-like property, are there some certain variational principles for metric mean dimensions of saturated sets? If not, can metric mean dimensions of saturated sets have the full metric mean dimension?
For (1), the authors in [YCZ24] showed the packing metric mean dimension of the set of generic points equals to the mean Rényi information dimension, however it holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula in infinite entropy systems (cf.[YCZ24, §5, Question,(1)]); for (2), noting the uniform separation property implies the system has finite topological entropy, so extending the work of [PS07] to the framework of infinite entropy systems encounters some nontrivial challenges. Only the assumption of $g$-almost product property, Yuan [Y23] investigated the Bowen metric mean dimension of $G_{K}$, and showed that for any $\{\mu_{1},\mu_{2},...,\mu_{k}\}\subset M_{f}(X)$ and any non-empty compact connected set $K$ contained in the convex closure ${\rm cov}\{\mu_{1},...,\mu_{k}\}$,
$\overline{\rm mdim}_{M}^{B}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Prior to the comparison of different types of specification-like property, we only focus on the systems with specification property and give a full answer to question (2). Actually, since the asymptotic behavior of orbits ${\rm orb}_{f}(x)=\{f^{n}(x):n\geq 0\}$ is a fundamental topic in ergodic theory, the following statements of main results involve more general (mixed) saturated sets. For simplicity, we write
@@
Saturated sets exhibit different dynamical behaviors in terms of different types of metric mean dimensions. There exist certain variational principles for Bowen and packing metric metric mean dimensions of saturated sets (Theorems 1.1 and 1.2), while the upper capacity metric mean dimension of saturated set has full metric mean dimension (Theorem 1.4).
@@
Let $(X,d,f)$ be a dynamical system with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, and $U\subset X$ is a nonempty open set, then
$\overline{\rm mdim}_{M}^{P}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{P}\left(f,G_{K}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\},d\right)$
$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where $\overline{\rm mdim}_{M}^{P}(f,Z,d)$ is the packing upper metric mean dimension of $Z$, $h_{\mu}(f,\xi)$ is the measure-theoretic entropy of $\mu$ w.r.t. $\xi$, and the infimum is taken over all finite Borel measurable partitions of $X$.
@@
Let $(X,d,f)$ be a dynamical system with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact connected set, and $U\subset X$ is a nonempty open set, then
$\overline{\rm mdim}_{M}^{B}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{B}\left(f,G_{K}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\},d\right)$
$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
@@
Maybe, more satisfactory results are exchanging the order of $\limsup_{\epsilon\to 0}$ and $\sup_{\mu\in K}$ (or $\inf_{\mu\in K}$), but until now the authors do not know whether there are some counter-examples of infinite entropy systems with specification property such that the following equalities are strict:
@@
Inspired by the work of [GS21, CPV24], we define two variants of mean Rényi information dimension of invariant measures. For any closed set $K$ of $M_{f}(X)$ and $\mu\in K$, by $M_{K}(\mu)$ we denote the space of sequences of Borel probability measures in $K$ which converge to $\mu$ in the weak-topology. We respectively define the (modified) lower and upper mean Rényi information dimensions of $\mu\in K$ as
$\underline{\rm mrid}_{\mu}(f,K,d):$ $=\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi),$
$\overline{\rm mrid}_{\mu}(f,K,d):$ $=\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
Then we have the following desired variational principles in terms of the modified lower and upper mean Rényi information dimensions.
@@
Let $(X,d,f)$ be a dynamical system with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, and $U\subset X$ is a nonempty open set, then
@@
Let $(X,d,f)$ be a dynamical system with specification property. Let $K\subseteq M_{f}(X)$ be a nonempty compact connected set, and $U\subset X$ be a nonempty open set, then
$\overline{\rm mdim}_{M}^{UC}\left(f,G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},d\right)=\overline{\rm mdim}_{M}^{UC}\left(f,G_{K},d\right)=\overline{\rm mdim}_{M}(f,X,d),$
where $\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ is the upper capacity metric mean dimension of $Z$ defined by separated sets.
As a direct consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case is $K$ being exactly a single point. The Corollary 1.5 presents the metric mean dimensions of generic points of any invariant measures for systems with specification property, which extends the result given in [YCZ24, Theorem 1.1] to all non-ergodic measures.
@@
Let $(X,d,f)$ be a dynamical system with specification property. Then for any $\mu\in M_{f}(X)$,
@@
Consequently, if $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure such that $\overline{\rm mdim}_{M}(f,X,d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi)$, then we have
@@
The “Consequently” part can be realized for some dynamical systems. For example, the full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$ with the compatible product metric $d(x,y)=\sum_{n\in\mathbb{Z}}\frac{||x_{n}-y_{n}||_{\mathbb{R}^{m}}}{2^{|n|}}$, and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ is product measure of the Lebesgue measure $\mathcal{L}$ on $[0,1]^{m}$. The concept of maximal metric mean dimension measure was proposed in [YCZ23a], yet for what kind of infinite entropy systems can have maximal metric mean dimension measure is still a difficult problem. See [LR24] for the existence of maximal metric mean dimension measure for typical conservative homeomorphisms. By Gutman- Śpiewak variational principle (see Theorem 2.4) and Corollary 1.5 we have following (abstract) criterion for maximal metric mean dimension measure for systems with specification property.
@@
Let $(X,d,f)$ be a dynamical system with specification property. An invariant measure $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure such that $\overline{\rm mdim}_{M}(f,X,d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi)$ if and only if $\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
@@
    There are plenty of infinite entropy systems with specification property, for instance, the usual full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$, and the examples presented in [BR23, Section IV] such that our main results can be applied to such systems.
@@
    In Theorem 1.3, we do not know whether the equalities hold for Bowen upper metric mean dimension and packing lower metric mean dimension. The other aforementioned theorems are valid for the corresponding lower metric mean dimensions by considering changing the $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.

### v3

@@
Abstract. Saturated sets and its reduced case, the set of generic points, constitute two significant types of fractal-like sets in multifractal analysis of dynamical systems. In the context of infinite entropy systems, we aim to provide some qualitative aspects of saturated sets and the set of generic points from both topological and measure-theoretic perspectives.
For systems with the specification property, we establish certain variational principles for saturated sets in terms of Bowen and packing metric mean dimensions, and show the upper capacity metric mean dimension of saturated sets has full metric mean dimension. As applications, we further present some qualitative aspects of the metric mean dimensions of level sets and the set of mean Li-Yorke pairs in infinite entropy systems.
@@ 1 Introduction
Let $(X,d,f)$ be a topological dynamical system (t.d.s. for short), where $(X,d)$ is a compact metric space with a metric $d$, and $f:X\to X$ is a homeomorphism. Let $M(X)$, $M_{f}(X)$, $M_{f}^{e}(X)$ denote the sets of Borel probability measures endowed with weak-topology, $f$-invariant Borel probability measures, and ergodic $f$-invariant Borel probability measures on $X$, respectively.
Quantitative analysis using some entropy-like quantities is of great significance in understanding the dynamics of abstract dynamical systems. Measure-theoretic entropy and topological entropy are two crucial invariants that helps us understand the topological complexity of dynamical systems, which are related by the classical variational principle:
@@
where $h_{top}(f)$ is the topological entropy of $X$, $h_{\mu}(f)$ is the measure-theoretic entropy of $\mu$, and the supremum ranges over all $f$-invariant measures of $X$. In 1973, resembling the definition of Hausdorff dimension Bowen [Bow73] introduced Bowen topological entropy for any subset of the phase space $X$. The pioneering work has directly contributed to the development of dimension theory in dynamical systems, especially serving as a powerful tool for studying the “size” of fractal-like sets in multifractal analysis. The multifractal analysis, with its origin dating back to Besicovitch, takes into account the decomposition of the whole phase space into smaller subsets consisting of points with the similar dynamical behaviors(or known as multifractal spectrum), and then, from a geometrical or topological perspective, describes the “size” of those subsets by using fractal-like dimensions and dimensional-like entropies. Such a procedure allows us to get more information about the dynamics and even recovers the partial dynamical behaviors of the system, a phenomenon known as multifractal rigidity.
We aim to study the multifractal analysis of saturated sets in infinite entropy systems. Let us mention some relevant results about the dimensional entropies of saturated sets to clarify the motivation. Given $x\in X$, by
$\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}\in M(X),$
where $\delta_{x}$ is the Dirac mass at $x$, we denote the $n$-th empirical measure of $x$. Then, in the sense of weak-topology, the limit-point set $V_{f}(x)$ of the sequence $\{\mathcal{E}_{n}(x)\}_{n\geq 1}$ is a nonempty compact connected subset of $M_{f}(X)$ (cf. [DGS76, Proposition 3.8]). Specially, when $V_{f}(x)=\{\mu\}$ for some invariant measure $\mu$, then $x$ is called a generic point of $\mu$. Denote by $G_{\mu}$ the set of all generic points of $\mu$. The Birkhoff’s ergodic theorem implies that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic. In 1973, Bowen [Bow73] proved that for any ergodic measure $\mu$, the Bowen topological entropy of $G_{\mu}$ equals $h_{\mu}(f)$, i.e.,
@@
For the case of non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$, and there even exist the examples showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$. Therefore, Bowen’s result fails for non-ergodic measures. It is natural to ask whether there are some proper conditions such that the Bowen’s equality still holds for non-ergodic measures.
A positive answer to the above question is due to Pfister and Sullivan [PS07], who provided it within a more general framework. A subset $K$ of $X$ is called a saturated set if $x\in K$ and $V_{f}(x)=V_{f}(y)$, then $y\in K$. More generally, given a nonempty closed connected subset $K\subset M_{f}(X)$, we have a saturated set by defining $G_{K}=\{x\in X:V_{f}(x)=K\}$ associated with $K$. For any topological dynamical system (t.d.s) possessing the specification property, Sigmund [Sig72, Sig74] firstly proved the existence of the saturated sets, and then was later extended to non-uniformly hyperbolic and non-uniformly expanding systems by several authors [LST13, TV17]. In 2007, Pfister and Sullivan [PS07] introduced the $g$-almost product property that is strictly weaker than specification property, which solely requires a point partially $\epsilon$-shadowing the given finite pieces of orbit-segments and the uniform separation property (e.g. expansive and more generally asymptotically $h$-expansive systems), and established the variational principle of Bowen topological entropy of nonempty compact connected set $K\subset M_{f}(X)$:
@@
Furthermore, under only the assumption of $g$-almost product property, they [PS07] showed that for all $\mu\in M_{f}(X)$,
$h_{top}^{B}(f,G_{\mu})=h_{\mu}(f),$
which extends Bowen’s remarkable result to non-ergodic measures. Consequently, the work of [Bow73, PS07] leads to an intensive study toward the estimation of different types of dimensional-like entropies of saturated sets. Under the specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] also obtained the variational principle of packing topological entropy of nonempty compact connected set $K\subset M_{f}(X)$,
$h_{top}^{P}(f,G_{K})=\sup\{h_{\mu}(f):\mu\in K\}.$
By weakening the specification property and considering systems with the $g$-almost product property, Hou, Tian and Zhang [HTZ23] not only demonstrated that the variational principle $h_{top}^{P}(f,G_{K}^{C})=\sup\{h_{\mu}(f):\mu\in K\}$ holds for (mixed) saturated sets with more complicated dynamics, but also showed that the upper capacity topological entropy of the saturated sets has full topological entropy, i.e.,
@@
where $G_{K}^{C}:=G_{K}\cap\{x\in X:C_{f}(X)\subset\omega_{f}(x)\},$ $\omega_{f}(x):=\bigcap_{n=0}^{\infty}\overline{\bigcup_{k=n}^{\infty}\{f^{k}x\}}$ is the $\omega$-limit set of $x$, and $C_{f}(X)$ is the measure center given by $C_{f}(X):=\overline{\cup_{\mu\in M_{f}(X)}{\rm supp}{(\mu)}}$, ${\rm supp}(\mu)$ is the supported set of $\mu$. Additionally, Huang, Tian and Wang [HTW19] showed that Pfister and Sullivan’s variational principle is still valid for transitive-saturated set $G_{K}^{T}:=G_{K}\cap{\rm Tran}$, where ${\rm Tran}$ means the set of transitive points. The generalizations of the aforementioned results can be found in [JCZ22, LLS17] for the scenarios of nonuniformly hyperbolic systems, and in [ZC18, DZZ23, Z18, RTZ23, ZRZ23] for the cases of the actions of amenable groups
Based on the possible values of topological entropy, dynamical systems can be can be classified into three types: zero entropy systems, finite positive entropy systems and infinite entropy systems. Infinite entropy systems typically imply that these systems may exhibit rather complicated topological complexity. As a result, the concept of topological entropy fails to offer more information about the dynamics of the system. In 2000, Lindenstrauss and Weiss [LW00] introduced a refined notion for topological entropy, called metric mean dimension, which is a dynamical analogue of box dimension to measure how fast the $\epsilon$-topological entropy diverges to $\infty$ as $\epsilon$ goes to 00. It has been shown that every system with finite entropy has a metric mean dimension of zero. See some applications of metric mean dimension in the analog compression of information theory [GS20], and in estimating the upper bound of mean dimension of some complex systems [Tsu18a, Tsu18b]. The notion of measure-theoretic counterpart of metric mean dimension, called upper and lower mean Rényi information dimensions, is due to Gutman and Śpiewak[GS21]. More precisely, the lower and upper mean Rényi information dimensions of $\mu\in M_{f}(X)$ are respectively defined by
@@
where the infimum ranges over the finite Borel measurable partition with diameter at most $\epsilon$. This measures how fast the Kolmogorov-Sinai $\epsilon$-entropy $\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ diverges to $\infty$ as $\epsilon\to 0$. Several variational principles regarding the metric mean dimension of $\mathbb{Z}$-actions can be found in the works [LW00, GS21, S22, YCZ23a, YCZ25]. Then, in the context of the infinite entropy systems, the aforementioned work suggests the following questions for saturated sets:
@@
    For ergodic measures, does the metric mean dimension of the set of generic points coincide with the mean Rényi information dimension?
@@
    For non-ergodic measures, under the assumption of specification-like property, are there certain variational principles for metric mean dimensions of saturated sets? If not, can metric mean dimensions of saturated sets have the full metric mean dimension?
For (1), the authors in [YCZ25] showed that the packing metric mean dimension of the set of generic points is equal to the mean Rényi information dimension. However, this equality holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula within the context of infinite entropy systems (cf.[YCZ25, §5, Question,(1)]); for (2), since the uniform separation property implies the system has finite topological entropy, extending the work of [PS07] to the framework of infinite entropy systems encounters some nontrivial challenges. Under the assumption of $g$-almost product property, for any non-empty compact connected set $K$ contained in the convex closure of finite invariant measures, Yuan [Y23] investigated the Bowen metric mean dimension of $G_{K}$.
Prior to the comparison of the different types of specification-like property, we concentrate solely on the systems possessing the specification property and provide a comprehensive answer to question (2). Actually, since the asymptotic behavior of orbits ${\rm orb}_{f}(x)=\{f^{n}(x):n\geq 0\}$ is a fundamental topic in ergodic theory, the subsequent statements of main results involve more general (mixed) saturated sets. For simplicity, we write
@@
Saturated sets exhibit different dynamical behaviors in terms of different types of metric mean dimensions. There are indeed certain variational principles for the Bowen and packing metric mean dimensions of saturated sets, as presented in Theorems 1.1, 1.2 and 1.3. Although the $\mu$-measure of $G_{\mu}$ is zero for every $\mu\in M_{f}(X)$, its topological complexity can be sufficiently large, that is, the upper capacity metric mean dimension of saturated sets carries the full metric mean dimension, as stated in Theorem 1.4.
@@
Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, and $U\subset X$ is a nonempty open set, then
$\overline{\rm mdim}_{M}^{P}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where $\overline{\rm mdim}_{M}^{P}(f,Z,d)$ is the packing upper metric mean dimension of $Z$, $h_{\mu}(f,\xi)$ is the measure-theoretic entropy of $\mu$ w.r.t. $\xi$.
@@
Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact connected set, and $U\subset X$ is a nonempty open set, then
$\overline{\rm mdim}_{M}^{B}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
@@
Perhaps, more satisfactory results would be to exchange the order of $\limsup_{\epsilon\to 0}$ and $\sup_{\mu\in K}$ (or $\inf_{\mu\in K}$). However, up to now, the authors remain unaware of whether there exist any counterexamples among infinite entropy systems with the specification property for which the following equalities hold strictly
@@
Inspired by the work of [GS21, CPV24], we define two variants of mean Rényi information dimension of invariant measures. Let $K$ be a closed set of $M_{f}(X)$. For any $\mu\in K$, we use $M_{K}(\mu)$ to denote the space of sequences of invariant measures within $K$ which converge to $\mu$ in the weak-topology. We define the (modified) lower and upper mean Rényi information dimensions of $\mu$ as
$\underline{\rm mrid}_{\mu}(f,K,d):$ $=\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\},$
$\overline{\rm mrid}_{\mu}(f,K,d):$ $=\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}$
respectively. Then this allows us to obtain the the variational principles in terms of the (modified) lower and upper mean Rényi information dimensions.
@@
Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, then
@@
Let $(X,d,f)$ be a t.d.s. with specification property. Let $K\subseteq M_{f}(X)$ be a nonempty compact connected set, and $U\subset X$ be a nonempty open set. Then
$\overline{\rm mdim}_{M}^{UC}\left(f,G_{K},d\right)=\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}(f,X,d),$
where $\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ is the upper capacity metric mean dimension of $Z$.
As an immediate consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case occurs when $K$ consists of exactly a single point. The Corollary 1.5 extends the result given in [YCZ25, Theorem 1.1] to all non-ergodic measures for systems with specification property.
@@
Let $(X,d,f)$ be a dynamical system with specification property. Then for every $\mu\in M_{f}(X)$,
@@
Consequently, if $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure in the sense of $\overline{\rm mdim}_{M}(f,X,d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$, then we have
@@
The “Consequently” part can be realized for some dynamical systems. For example, the full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$ with the compatible product metric $d(x,y)=\sum_{n\in\mathbb{Z}}\frac{||x_{n}-y_{n}||_{\mathbb{R}^{m}}}{2^{|n|}}$, and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ is product measure of the Lebesgue measure $\mathcal{L}$ on $[0,1]^{m}$.
The concept of maximal metric mean dimension measure was proposed in [YCZ23a]. However, determining what types of infinite entropy systems can possess a maximal metric mean dimension measure remains a challenging task. See [LR24] for the existence of maximal metric mean dimension measure for typical conservative homeomorphisms. By Gutman- Śpiewak variational principle (as presented in Theorem 2.4) and Corollary 1.5, we have an abstract criterion for maximal metric mean dimension measure for identifying maximal metric mean dimension measures in systems with the specification property.
@@
Let $(X,d,f)$ be a t.d.s. with specification property. An invariant measure $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure if and only if $\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
@@
    There are plenty of infinite entropy systems with specification property, for instance, the usual full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$, and the examples presented in [BR23, Section IV], and our main results can be applied to these systems.
@@
    In Theorem 1.3, we do not know whether the equalities hold for Bowen upper metric mean dimension and packing lower metric mean dimension. The other theorems mentioned above are valid for the corresponding lower metric mean dimensions by considering changing the $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.